# Deep learning empowered sensor fusion to improve infant movement classification

### v1

@@ Acknowledgments
The authors thank all families for their participation in this close-meshed prospective study. The initial idea of this study was conceptualized together with the “father of GMA”, Heinz Prechtl, and one of the leading experts in spontaneous infant movement assessment, Christa Einspieler. The transition from a Gestalt to an AI-based GMA would not have been possible without their guidance and skepticism on our initial ideas which was essential to develop the suggested methodology. Special thanks to the team members of the Marschik-Labs in Graz and Göttingn (iDN–interdisciplinary Developmental Neuroscience and SEE–Systemic Ethology and Developmental Science), who were involved in recruitment, accompanying families and their babies, data acquisition, data curation and pre-processing, and technical support: Gunter Vogrinec, Magdalena Krieber-Tomantschger, Iris Tomantschger, Laura Langmann, Claudia Zitta, Dr. Robert Peharz, Dr. Florian Pokorny, Dr. Simon Reich, Lennart Jahn, and Sarah Flügge. Special thanks also to our interdisciplinary international network of collaborators for discussing this study with us and for refining our views and ideas.
@@ 1 Pre-processing of the video data
where $k=1\dots 15$ is the key point index, and $f=1\dots n$ ($n=250$) is the frame index. $x^{p_{10,11}}_{f}$ and $y^{p_{10,11}}_{f}$ correspond to the $X$ and $Y$ coordinates of the hip key points (see Figure 2c).
@@ 2 Pre-processing of the pressure mat data
We first cropped the area of size 29$\times$26 ([1:29, 4:29]) of the original grid size (see red rectangle in Figure 2b), since the sensor values outside this area were 0 in most of the cases. The cropped area contained 754 pressure sensor values. Generally, only two areas were strongly activated on the pressure mat (see Figure 2a). The activation at the top corresponds to the infants’ shoulders and/or head, whereas activation at the bottom corresponds to the infant’s hips. Thus, we split the cropped grid area of size 29$\times$26 into two parts, 12$\times$26 (top part) and 17$\times$26 (bottom part), and tracked the center of pressure (CoP) in these two areas.

### v2

@@ Acknowledgments
The authors thank all families for their participation in this close-meshed prospective study. The initial idea of this study was conceptualized together with the “father of GMA”, Heinz Prechtl, and one of the leading experts in spontaneous infant movement assessment, Christa Einspieler. The transition from a Gestalt to an AI-based GMA would not have been possible without their guidance and skepticism on our initial ideas which was essential to develop the suggested methodology. Special thanks to the team members of the Marschik-Labs in Graz and Göttingn (iDN–interdisciplinary Developmental Neuroscience and SEE–Systemic Ethology and Developmental Science), who were involved in recruitment, accompanying families and their babies, data acquisition, data curation and pre-processing, and technical support: Gunter Vogrinec, Magdalena Krieber-Tomantschger, Iris Tomantschger, Laura Langmann, Claudia Zitta, Dr. Robert Peharz, Dr. Florian Pokorny, and Dr. Simon Reich. Special thanks also to our interdisciplinary international network of collaborators for discussing this study with us and for refining our views and ideas.
@@ 1 Pre-processing of the video data
where $k=1\dots 15$ is the key point index, and $f=1\dots n$ ($n=250$) is the frame index. $x^{p_{10,11}}_{f}$ and $y^{p_{10,11}}_{f}$ correspond to the $X$ and $Y$ coordinates of the hip key points (see Figure 2a).
@@ 2 Pre-processing of the pressure mat data
We first cropped the area of size 29$\times$26 ([1:29, 4:29]) of the original grid size (see red rectangle in Figure 2b), since the sensor values outside this area were 0 in most of the cases. The cropped area contained 754 pressure sensor values. Generally, only two areas were strongly activated on the pressure mat (see Figure 2b). The activation at the top corresponds to the infants’ shoulders and/or head, whereas activation at the bottom corresponds to the infant’s hips. Thus, we split the cropped grid area of size 29$\times$26 into two parts, 12$\times$26 (top part) and 17$\times$26 (bottom part), and tracked the center of pressure (CoP) in these two areas.